\def\cprime{$'$}
\begin{thebibliography}{10}

\bibitem{Abramowitz-Stegun-1964}
M.~Abramowitz and I.~A. Stegun.
\newblock {\em Handbook of mathematical functions with formulas, graphs, and
  mathematical tables}.
\newblock U. S. Government Printing Office, Washington, DC, 1964.

\bibitem{Albeverio-Brzezniak-Wu-2010}
S.~Albeverio, Z.~Brze\'{z}niak, and J.~Wu.
\newblock Existence of global solutions and invariant measures for stochastic
  differential equations driven by {P}oisson type noise with non-{L}ipschitz
  coefficients.
\newblock {\em J. Math. Anal. Appl.}, 371(1):309--322, 2010.

\bibitem{Arapostathis-Borkar-Ghosh_Book-2012}
A.~Arapostathis, V.~S. Borkar, and M.~K. Ghosh.
\newblock {\em Ergodic control of diffusion processes}.
\newblock Cambridge University Press, Cambridge, 2012.

\bibitem{Arapostathis-Pang-Sandric-2019}
A.~Arapostathis, G.~Pang, and N.~Sandri\'c.
\newblock Ergodicity of a {L}\'evy--driven {SDE} arising from multiclass
  many-server queues.
\newblock {\em Ann. Appl. Probab.}, 29(2):1070--1126, 2019.

\bibitem{Arisawa-2009}
M.~Arisawa.
\newblock Homogenization of a class of integro-differential equations with
  {L}\'{e}vy operators.
\newblock {\em Comm. Partial Differential Equations}, 34(7-9):617--624, 2009.

\bibitem{Bass-Cranston-1986}
R.~F. Bass and M.~Cranston.
\newblock The {M}alliavin calculus for pure jump processes and applications to
  local time.
\newblock {\em Ann. Probab.}, 14(2):490--532, 1986.

\bibitem{Bhattacharya-1978}
R.~N. Bhattacharya.
\newblock Criteria for recurrence and existence of invariant measures for
  multidimensional diffusions.
\newblock {\em Ann. Probab.}, 6(4):541--553, 1978.

\bibitem{Bolley-Gentil-Guillin-2012}
F.~Bolley, I.~Gentil, and A.~Guillin.
\newblock Convergence to equilibrium in {W}asserstein distance for
  {F}okker-{P}lanck equations.
\newblock {\em J. Funct. Anal.}, 263(8):2430--2457, 2012.

\bibitem{Butkovsky-2014}
O.~Butkovsky.
\newblock Subgeometric rates of convergence of {M}arkov processes in the
  {W}asserstein metric.
\newblock {\em Ann. Appl. Probab.}, 24(2):526--552, 2014.

\bibitem{Chen-Chen-Wang-2020}
X.~Chen, Z.-Q. Chen, and J.~Wang.
\newblock Heat kernel for non-local operators with variable order.
\newblock {\em Stochastic Process. Appl.}, 130(6):3574--3647, 2020.

\bibitem{Chen-Hu-Xie-Zhang-2017}
Z.-Q. Chen, E.~Hu, L.~Xie, and X.~Zhang.
\newblock Heat kernels for non-symmetric diffusion operators with jumps.
\newblock {\em J. Differential Equations}, 263(10):6576--6634, 2017.

\bibitem{Chen-Zhang-2016}
Z.-Q. Chen and X.~Zhang.
\newblock Heat kernels and analyticity of non-symmetric jump diffusion
  semigroups.
\newblock {\em Probab. Theory Related Fields}, 165(1-2):267--312, 2016.

\bibitem{Chen-Zhang-2018}
Z.-Q. Chen and X.~Zhang.
\newblock Heat kernels for time-dependent non-symmetric stable-like operators.
\newblock {\em J. Math. Anal. Appl.}, 465(1):1--21, 2018.

\bibitem{Deng-Schilling-Song-2017}
C.-S. Deng, R.~L. Schilling, and Y.-H. Song.
\newblock Subgeometric rates of convergence for {M}arkov processes under
  subordination.
\newblock {\em Adv. in Appl. Probab.}, 49(1):162--181, 2017.

\bibitem{Deng-Schilling-Song-Erratum-2018}
C.-S. Deng, R.~L. Schilling, and Y.-H. Song.
\newblock Correction: ``{S}ubgeometric rates of convergence for {M}arkov
  processes under subordination'' [ {MR}3631220].
\newblock {\em Adv. in Appl. Probab.}, 50(3):1005, 2018.

\bibitem{Douc-Fort-Guilin-2009}
R.~Douc, G.~Fort, and A.~Guillin.
\newblock Subgeometric rates of convergence of {$f$}-ergodic strong {M}arkov
  processes.
\newblock {\em Stochastic Process. Appl.}, 119(3):897--923, 2009.

\bibitem{Durmus-Fort-Moulines-Soulier-2004}
R.~Douc, G.~Fort, E.~Moulines, and P.~Soulier.
\newblock Practical drift conditions for subgeometric rates of convergence.
\newblock {\em Ann. Appl. Probab.}, 14(3):1353--1377, 2004.

\bibitem{Douc-Moulines-Priouret-Soulier_Book-2018}
R.~Douc, E.~Moulines, P.~Priouret, and P.~Soulier.
\newblock {\em Markov chains}.
\newblock Springer, Cham, 2018.

\bibitem{Down-Meyn-Tweedie-1995}
D.~Down, S.~P. Meyn, and R.~L. Tweedie.
\newblock Exponential and uniform ergodicity of {M}arkov processes.
\newblock {\em Ann. Probab.}, 23(4):1671--1691, 1995.

\bibitem{Durmus-Fort-Moulines-2016}
A.~Durmus, G.~Fort, and E.~Moulines.
\newblock Subgeometric rates of convergence in {W}asserstein distance for
  {M}arkov chains.
\newblock {\em Ann. Inst. Henri Poincar\'{e} Probab. Stat.}, 52(4):1799--1822,
  2016.

\bibitem{Durrett-Book-1996}
R.~Durrett.
\newblock {\em Stochastic calculus}.
\newblock CRC Press, Boca Raton, FL, 1996.

\bibitem{Eberle-2011}
A.~Eberle.
\newblock Reflection coupling and {W}asserstein contractivity without
  convexity.
\newblock {\em C. R. Math. Acad. Sci. Paris}, 349(19-20):1101--1104, 2011.

\bibitem{Eberle-2015}
A.~Eberle.
\newblock Reflection couplings and contraction rates for diffusions.
\newblock {\em Probab. Theory Related Fields}, 166(3-4):851--886, 2016.

\bibitem{Fort-Moulines-2003}
G.~Fort and E.~Moulines.
\newblock Polynomial ergodicity of {M}arkov transition kernels.
\newblock {\em Stochastic Process. Appl.}, 103(1):57--99, 2003.

\bibitem{Fort-Roberts-2005}
G.~Fort and G.~O. Roberts.
\newblock Subgeometric ergodicity of strong {M}arkov processes.
\newblock {\em Ann. Appl. Probab.}, 15(2):1565--1589, 2005.

\bibitem{Friedman-Book-1975}
A.~Friedman.
\newblock {\em Stochastic differential equations and applications. {V}ol. 1}.
\newblock Academic Press [Harcourt Brace Jovanovich, Publishers], New
  York-London, 1975.

\bibitem{Grzywny-Szczypkowski-2019}
T.~Grzywny and K.~Szczypkowski.
\newblock Heat kernels of non-symmetric {L}\'{e}vy-type operators.
\newblock {\em J. Differential Equations}, 267(10):6004--6064, 2019.

\bibitem{Hairer-Mattingly-Scheutzow-2011}
M.~Hairer, J.~C. Mattingly, and M.~Scheutzow.
\newblock Asymptotic coupling and a general form of {H}arris' theorem with
  applications to stochastic delay equations.
\newblock {\em Probab. Theory Related Fields}, 149(1-2):223--259, 2011.

\bibitem{Ishikawa-2001}
Y.~Ishikawa.
\newblock Density estimate in small time for jump processes with singular
  {L}\'evy measures.
\newblock {\em Tohoku Math. J. (2)}, 53(2):183--202, 2001.

\bibitem{Kevei-2018}
P.~Kevei.
\newblock Ergodic properties of generalized {O}rnstein-{U}hlenbeck processes.
\newblock {\em Stochastic Process. Appl.}, 128(1):156--181, 2018.

\bibitem{Khasminskii-1960}
R.~Z. Khasminskii.
\newblock Ergodic properties of recurrent diffusion processes and stabilization
  of the solution of the {C}auchy problem for parabolic equations.
\newblock {\em Teor. Verojatnost. i Primenen.}, 5:196--214, 1960.

\bibitem{Kim-Lee-2019}
P.~Kim and J.~Lee.
\newblock Heat kernels of non-symmetric jump processes with exponentially
  decaying jumping kernel.
\newblock {\em Stochastic Process. Appl.}, 129(6):2130--2173, 2019.

\bibitem{Kim-Song-Vondracek-2018}
P.~Kim, R.~Song, and Z.~Vondra\v{c}ek.
\newblock Heat kernels of non-symmetric jump processes: beyond the stable case.
\newblock {\em Potential Anal.}, 49(1):37--90, 2018.

\bibitem{Knopova-Kulik-2014}
V.~Knopova and A.~M. Kulik.
\newblock Parametrix construction of the transition probability density of the
  solution to an {SDE} driven by {$\alpha$}-stable noise.
\newblock {\em Ann. Inst. Henri Poincar\'{e} Probab. Stat.}, 54(1):100--140,
  2018.

\bibitem{Knopova-Schilling-2012}
V.~P. Knopova and R.~L. Schilling.
\newblock Transition density estimates for a class of {L}\'evy and
  {L}\'evy-type processes.
\newblock {\em J. Theoret. Probab.}, 25(1):144--170, 2012.

\bibitem{Knopova-Schilling-2013}
V.~P. Knopova and R.~L. Schilling.
\newblock A note on the existence of transition probability densities of
  {L}\'evy processes.
\newblock {\em Forum Math.}, 25(1):125--149, 2013.

\bibitem{Kolokoltsov-2000}
V.~N. Kolokoltsov.
\newblock Symmetric stable laws and stable-like jump-diffusions.
\newblock {\em Proc. Lond. Math. Soc. (3)}, 80(3):725--768, 2000.

\bibitem{Kolokoltsov-Book-2011}
V.~N. Kolokoltsov.
\newblock {\em Markov processes, semigroups and generators}, volume~38.
\newblock Walter de Gruyter \& Co., Berlin, 2011.

\bibitem{Kulik-2009}
A.~M. Kulik.
\newblock Exponential ergodicity of the solutions to {SDE}'s with a jump noise.
\newblock {\em Stochastic Process. Appl.}, 119(2):602--632, 2009.

\bibitem{Kulik-Book-2015}
A.~M. Kulik.
\newblock {\em Introduction to ergodic rates for {M}arkov chains and
  processes}.
\newblock Potsdam University Press, Potsdam, 2015.

\bibitem{Kulik-Book-2018}
A.~M. Kulik.
\newblock {\em Ergodic behavior of {M}arkov processes}.
\newblock De Gruyter, Berlin, 2018.

\bibitem{Kwon-Lee-1999}
Y.~Kwon and C.~Lee.
\newblock Strong {F}eller property and irreducibility of diffusions with jumps.
\newblock {\em Stochastics Stochastics Rep.}, 67(1-2):147--157, 1999.

\bibitem{Lazic-Sandric-2021}
P.~Lazi\'{c} and N.~Sandri\'{c}.
\newblock On sub-geometric ergodicity of diffusion processes.
\newblock {\em Bernoulli}, 27(1):348--380, 2021.

\bibitem{Lazic-Sandric-2022}
P.~Lazi\'{c} and N.~Sandri\'{c}.
\newblock On subgeometric ergodicity of regime-switching diffusion processes.
\newblock {\em Nonlinear Anal. Hybrid Syst.}, 46:Paper No. 101262, 24, 2022.

\bibitem{Luo-Wang-2016}
D.~Luo and J.~Wang.
\newblock Exponential convergence in {$L^p$}-{W}asserstein distance for
  diffusion processes without uniformly dissipative drift.
\newblock {\em Math. Nachr.}, 289(14-15):1909--1926, 2016.

\bibitem{Majka-2017}
M.~B. Majka.
\newblock Coupling and exponential ergodicity for stochastic differential
  equations driven by {L}\'{e}vy processes.
\newblock {\em Stochastic Process. Appl.}, 127(12):4083--4125, 2017.

\bibitem{Majka-2020}
M.~B. Majka.
\newblock A note on existence of global solutions and invariant measures for
  jump {SDE}s with locally one-sided {L}ipschitz drift.
\newblock {\em Probab. Math. Statist.}, 40(1):37--55, 2020.

\bibitem{Mao-2002}
Y.-H. Mao.
\newblock Strong ergodicity for {M}arkov processes by coupling methods.
\newblock {\em J. Appl. Probab.}, 39(4):839--852, 2002.

\bibitem{Mao-2006}
Y.-H. Mao.
\newblock Convergence rates in strong ergodicity for {M}arkov processes.
\newblock {\em Stochastic Process. Appl.}, 116(12):1964--1976, 2006.

\bibitem{Masuda-2007}
H.~Masuda.
\newblock Ergodicity and exponential {$\beta$}-mixing bounds for
  multidimensional diffusions with jumps.
\newblock {\em Stochastic Process. Appl.}, 117(1):35--56, 2007.

\bibitem{Masuda-Erratum-2009}
H.~Masuda.
\newblock Erratum to: ``{E}rgodicity and exponential {$\beta$}-mixing bound for
  multidimensional diffusions with jumps'' [{S}tochastic {P}rocess. {A}ppl. 117
  (2007) 35--56].
\newblock {\em Stochastic Process. Appl.}, 119(2):676--678, 2009.

\bibitem{Meyn-Tweedie-Book-2009}
S.~Meyn and R.~L. Tweedie.
\newblock {\em Markov {C}hains and {S}tochastic {S}tability}.
\newblock Cambridge University Press, Cambridge, second edition, 2009.

\bibitem{Meyn-Tweedie-AdvAP-II-1993}
S.~P. Meyn and R.~L. Tweedie.
\newblock Stability of {M}arkovian processes. {II}. {C}ontinuous-time processes
  and sampled chains.
\newblock {\em Adv. in Appl. Probab.}, 25(3):487--517, 1993.

\bibitem{Meyn-Tweedie-AdvAP-III-1993}
S.~P. Meyn and R.~L. Tweedie.
\newblock Stability of {M}arkovian processes. {III}. {F}oster-{L}yapunov
  criteria for continuous-time processes.
\newblock {\em Adv. in Appl. Probab.}, 25(3):518--548, 1993.

\bibitem{Pang-Sandric-2016}
G.~Pang and N.~Sandri\'{c}.
\newblock Ergodicity and fluctuations of a fluid particle driven by diffusions
  with jumps.
\newblock {\em Commun. Math. Sci.}, 2(14):327--362, 2016.

\bibitem{Picard-1996}
J.~Picard.
\newblock On the existence of smooth densities for jump processes.
\newblock {\em Probab. Theory Related Fields}, 105(4):481--511, 1996.

\bibitem{Picard-Erratum-2010}
J.~Picard.
\newblock Erratum to: {O}n the existence of smooth densities for jump
  processes.
\newblock {\em Probab. Theory Related Fields}, 147(3-4):711--713, 2010.

\bibitem{Prevot-Rockner-Book-2007}
C.~Pr\'ev\^ot and M.~R\"ockner.
\newblock {\em A concise course on stochastic partial differential equations}.
\newblock Springer, Berlin, 2007.

\bibitem{Rogers-Williams-Book-II-2000}
L.~C.~G. Rogers and D.~Williams.
\newblock {\em Diffusions, {M}arkov processes, and martingales. {V}ol. 2}.
\newblock Cambridge Mathematical Library. Cambridge University Press,
  Cambridge, 2000.

\bibitem{Sandric-ESAIM-2016}
N.~Sandri\'c.
\newblock Ergodicity of {L}\'evy-type processes.
\newblock {\em ESAIM-Prob.-Stat.}, 20:154--177, 2016.

\bibitem{Sandric-TAMS-2016}
N.~Sandri\'{c}.
\newblock Long-time behavior for a class of {F}eller processes.
\newblock {\em Trans. Am. Math. Soc.}, 3(368):1871--1910, 2016.

\bibitem{Sandric-Valentic-Wang-2021}
N.~Sandri\'{c}, I.~Valenti\'{c}, and J.~Wang.
\newblock Periodic homogenization of a {L}\'{e}vy-type process with small
  jumps.
\newblock {\em J. Evol. Equ.}, 21(1):771--803, 2021.

\bibitem{Sato-Book-1999}
K.~Sato.
\newblock {\em L\'evy processes and infinitely divisible distributions},
  volume~68.
\newblock Cambridge University Press, Cambridge, 1999.

\bibitem{Shao-Xi-2013}
J.~Shao and F.~Xi.
\newblock Strong ergodicity of the regime-switching diffusion processes.
\newblock {\em Stochastic Process. Appl.}, 123(11):3903--3918, 2013.

\bibitem{Stramer-Tweedie-1997}
O.~Stramer and R.~L. Tweedie.
\newblock Existence and stability of weak solutions to stochastic differential
  equations with non-smooth coefficients.
\newblock {\em Statist. Sinica}, 7(3):577--593, 1997.

\bibitem{Stroock-1975}
D.~W. Stroock.
\newblock Diffusion processes associated with {L}\'{e}vy generators.
\newblock {\em Z. Wahrscheinlichkeitstheorie und Verw. Gebiete},
  32(3):209--244, 1975.

\bibitem{Tuominen-Tweedie-1994}
P.~Tuominen and R.~L. Tweedie.
\newblock Subgeometric rates of convergence of {$f$}-ergodic {M}arkov chains.
\newblock {\em Adv. in Appl. Probab.}, 26(3):775--798, 1994.

\bibitem{Tweedie-1994}
R.~L. Tweedie.
\newblock Topological conditions enabling use of {H}arris methods in discrete
  and continuous time.
\newblock {\em Acta Appl. Math.}, 34(1-2):175--188, 1994.

\bibitem{Veretennikov-1997}
A.~Y. Veretennikov.
\newblock On polynomial mixing bounds for stochastic differential equations.
\newblock {\em Stochastic Process. Appl.}, 70, 1997.

\bibitem{Veretennikov-2000}
A.~Y. Veretennikov.
\newblock On polynomial mixing and the rate of convergence for stochastic
  differential and difference equations.
\newblock {\em Teor. Veroyatnost. i Primenen.}, 44(2):312--327, 1999.

\bibitem{Renesse-Sturm-2005}
M.~von Renesse and K.~Sturm.
\newblock Transport inequalities, gradient estimates, entropy, and {R}icci
  curvature.
\newblock {\em Comm. Pure Appl. Math.}, 58(7):923--940, 2005.

\bibitem{FYWang-2011}
F.-Y. Wang.
\newblock Coupling for {O}rnstein-{U}hlenbeck processes with jumps.
\newblock {\em Bernoulli}, 17(4):1136--1158, 2011.

\bibitem{Wang-2008}
J.~Wang.
\newblock Criteria for ergodicity of {L}\'evy type operators in dimension one.
\newblock {\em Stochastic Process. Appl.}, 118(10):1909--1928, 2008.

\bibitem{Wang-2011}
J.~Wang.
\newblock Stability of {M}arkov processes generated by {L}\'evy-type operators.
\newblock {\em Chinese Ann. Math. Ser. A}, 32(1):33--50, 2011.

\bibitem{Wang-2016}
J.~Wang.
\newblock {$L^p$}-{W}asserstein distance for stochastic differential equations
  driven by {L}\'{e}vy processes.
\newblock {\em Bernoulli}, 22(3):1598--1616, 2016.

\bibitem{Wee-1999}
I.~Wee.
\newblock Stability for multidimensional jump-diffusion processes.
\newblock {\em Stochastic Process. Appl.}, 80(2):193--209, 1999.

\bibitem{Xi-Zhu-2019}
F.~Xi and C.~Zhu.
\newblock Jump type stochastic differential equations with non-{L}ipschitz
  coefficients: non-confluence, {F}eller and strong {F}eller properties, and
  exponential ergodicity.
\newblock {\em J. Differential Equations}, 266(8):4668--4711, 2019.

\end{thebibliography}

\documentclass[12pt,a4paper, oneside, reqno]{amsart}
\usepackage[foot]{amsaddr} 
\usepackage{mathtools}
\usepackage{amsmath, nicefrac, amsthm, verbatim, amsfonts, mathtools, amssymb, upgreek, xcolor, bbm}
\usepackage{amsmath, nicefrac, amsthm, verbatim, amsfonts, amssymb, xcolor, bbm}
\usepackage{graphics, xspace, enumerate}
\usepackage{enumitem}
\usepackage{stix}
\usepackage[top=30mm,bottom=30mm,left=25mm,right=25mm]{geometry}

\usepackage[colorinlistoftodos,  textsize=tiny]{todonotes} 					 
\presetkeys{todonotes}{fancyline, color=blue!30}{}

\usepackage[bb=boondox]{mathalfa}
\usepackage{graphicx}
\usepackage[colorlinks=true,citecolor=red,urlcolor=blue,linkcolor=red,bookmarksopen=true,unicode=true,pdffitwindow=false]{hyperref}
\usepackage[english]{babel}
\usepackage[languagenames,fixlanguage]{babelbib}
\hypersetup{pdfauthor={}}
\hypersetup{pdftitle={A note on the uniform ergodicity of diffusion processes}}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{proposition}[theorem]{Proposition}
\theoremstyle{definition}

\newtheorem{example}[theorem]{Example}

\newcommand {\Prob} {\ensuremath{\mathbb{P}}}
\newcommand {\R} {\ensuremath{\mathbb{R}}}
\newcommand {\ZZ} {\ensuremath{\mathbb{Z}}}
\newcommand {\N} {\ensuremath{\mathbb{N}}}

\newcommand{\df}{\coloneqq}

\newcommand{\E}{\mathrm{e}}

\newcommand{\X}{\mathrm{X}}

\newcommand{\calC}{\mathcal{C}}

\newcommand{\D}{\mathrm{d}}

\numberwithin{equation}{section}

\title{A note on the uniform ergodicity of diffusion processes}

\author[N.\ Sandri\'{c}]{Nikola Sandri\'{c}}
\address[Nikola\ Sandri\'{c}]{Department of Mathematics\\University of Zagreb\\ Zagreb\\Croatia}
\email{nsandric@math.hr}

\makeatletter
\@namedef{subjclassname@2020}{%
	\textup{2020} Mathematics Subject Classification}
\makeatother

\subjclass[2020]{
60J25, 60J60, 60G17
}
\keywords{diffusion process, uniform ergodicity, total variation distance}

\begin{document}
\allowdisplaybreaks[4]

\begin{abstract}
 In this note, we discuss the uniform ergodicity of a diffusion process given by an Itô stochastic differential equation. We present an integral condition in terms of the drift and diffusion coefficients that ensures the uniform ergodicity of the corresponding transition kernel with respect to the total variation distance. Applications of the obtained results to a class of subordinate diffusion processes are also presented.
\end{abstract}

\maketitle

\section{Introduction}

One of the classical directions in the analysis of Markov processes centers around their ergodicity properties. In this note, we study the uniform ergodicity of a diffusion process given by  
\begin{equation}\begin{aligned} \label{eq1}\D X(x,t)&=b(X(x,t)) +\sigma(X(x,t))\D B(t)\\X(x,0)&=x\in\R^{d}\end{aligned}\end{equation}
with respect to the total variation distance.
Here, $\{B(t)\}_{t\ge0}$ stands for a standard $n$-dimensional Brownian motion (defined on a stochastic basis $(\Omega,\mathcal{F},\{\mathcal{F}(t)\}_{t\ge0},\mathbb{P})$ satisfying the usual conditions), and the coefficients $b:\R^d\to\R^d$ and $\sigma:\R^d\to\R^{d\times n}$ satisfy the following conditions: 

\medskip

\begin{description}
	\item[(A1)] For any $r>0$,
	$$\sup_{x\in B_r(0)}(\rvert b(x)\rvert+\lVert\sigma(x)\lVert_{{\rm HS}})<\infty.$$
	\item[(A2)] For any $r>0$, there exists $\Gamma_r>0$ such that for all $x,y\in B_r(0)$,  $$2\langle x-y,b(x)-b(y)\rangle+\lVert\sigma(x)-\sigma(y)\lVert_{{\rm HS}}^2\le \Gamma_r|x-y|^{2}.$$
	\item[(A3)] There exists $\Gamma>0$ such that for all $x\in\R^d$,  $$2\langle x,b(x)\rangle+\lVert\sigma(x)\lVert_{{\rm HS}}^{2}\le\Gamma(1+|x|^2).$$
	\item[(A4)] There exist  $x_0\in\R^d$ and $r_0,\alpha,D,\delta>0$,  such that for all   $x,y\in B_{r_0}(x_0)$,
	$$|b(x)-b(y)|+\lVert \sigma(x)\sigma(x)^T-\sigma(y)\sigma(y)^T\rVert_{{\rm HS}}\le D|x-y|^{\alpha}\quad\textrm{and}\quad \langle \sigma(x)^Ty,\sigma(x)^Ty\rangle\ge \delta|y|^{2}.$$
		\item[(A5)] For $x_0\in\R^d$ and $r_0>0$ from (A4),  $\sigma(x)\sigma(x)^T$ is positive definite on $B^c_{r_0}(x_0)$.
\end{description}

\medskip

\noindent Here,  $B_r(x)$ denotes the open ball with radius $r>0$ around $x\in\R^d$, and  $\lVert M\lVert_{{\rm HS}}^2:={\rm Tr}\,MM^T$ is the  Hilbert-Schmidt norm of a real  matrix $M.$

\subsection{Structural properties of the model} \label{ss1}
Under (A1)-(A3), for any $x\in\R^d$, the stochastic differential equation (SDE) in \eqref{eq1} admits a unique strong non-explosive solution $\{X(x,t)\}_{t\ge0}$ which is a strong Markov process with continuous sample paths and  transition kernel $p(t,x,\D y)=\Prob(X(x,t)\in \D y)$, $t\ge0$, $x\in\R^d$. 
See  \cite[Theorems 5.4.1, 5.4.5 and 5.4.6]{Durrett-Book-1996} and \cite[Theorem 3.1.1]{Prevot-Rockner-Book-2007} for more details. In the context of Markov processes, it is natural that the underlying probability measure depends on the initial distribution of the process. Using standard arguments (Kolmogorov extension theorem), it is well known that for each $x\in\R^d$ the above-defined transition kernel defines a unique probability measure $\mathbb{P}^x$ on the canonical (sample-path)  space such that the projection process, denoted by $\{X(t)\}_{t\ge0}$, is a strong Markov process (with respect to the completion of the corresponding natural filtration), it has continuous  sample paths, 
and  the same finite-dimensional distributions (with respect to $\mathbb{P}^x$) as $\{X(x,t)\}_{t\ge0}$ (with respect to $\mathbb{P}$). Since we are interested in distributional properties of the solution to   \eqref{eq1}, we focus on $\{X(t)\}_{t\ge0}$ rather than  $\{X(x,t)\}_{t\ge0}$ in the sequel. Furthermore, according to \cite[Lemma 2.5]{Majka-2020}, for any $t>0$ and continuous and bounded $f:\R^d\to\R$ the mapping $$x\mapsto\int_{\R^d}f(y)p(t,x,\D y)$$ is also continuous and bounded, i.e., $\{X(t)\}_{t\ge0}$ satisfies the so-called $\calC_b$-Feller property.   
In particular, this automatically implies that $\{X(t)\}_{t\ge0}$ is a strong Markov process with respect to the right-continuous and completed version of the  natural filtration.
Finally, in \cite[Theorem V.21.1]{Rogers-Williams-Book-II-2000} it is shown that
the class $\mathcal{C}^2(\R^d)$ is contained in the domain of the extended generator $\mathcal{G}$ of $\{X(t)\}_{t\ge0}$, and $\mathcal{G}$ restricted to this class has the following form:
\begin{equation}\label{gen}\mathcal{G}f(x)=\langle b(x),\nabla f(x)\rangle+\frac{1}{2}{\rm Tr}\,\sigma(x)\sigma(x)^T\nabla^2f(x).\end{equation}

\subsection{Main result} \label{ss2} Before stating the main result of this article, we first recall definitions of ergodicity. 

\medskip

\begin{enumerate}
\item [(i)]  A
(not necessarily finite) measure $\uppi(\D x)$ on $\mathcal{B}(\R^d)$ is called invariant for
$\{X(t)\}_{t\ge0}$ if
$$\int_{\R^d}p(t,x,\D y)\uppi(\D x)
= \uppi(\D y),\qquad t>0.$$
	\item [(ii)] The process $\{X(t)\}_{t\ge0}$  is said to be ergodic
	if it possesses an invariant probability 
	measure $\uppi(\D x)$  such that 
	\begin{equation*}
	\lim_{t\to\infty}\lVert p(t,x,\D {y})
	-\uppi(\D {y})\rVert_{{\rm TV}} =0,\qquad x \in \R^d.
	\end{equation*}
		\item [(iii)] The process $\{X(t)\}_{t\ge0}$  is said to be uniformly ergodic 
	if it possesses an invariant probability 
	measure $\uppi(\D x)$  such that 
	\begin{equation*}
	\lim_{t\to\infty}\sup_{x\in \R^d}\lVert p(t,x,\D {y})
	-\uppi(\D {y})\rVert_{{\rm TV}} =0.
	\end{equation*}
\end{enumerate}

\medskip

\noindent Here, $\lVert\upmu\rVert_{{\rm TV}}\df\sup_{B\in\mathcal{B}(\R^d)}|\upmu(B)|$ is the total variation norm  of a signed measure $\upmu(\D x)$ (on $\mathcal{B}(\R^d)$). Observe that: 

\medskip

\begin{enumerate}
	\item [(a)] (Uniform) ergodicity immediately implies the uniqueness of $\uppi(\D x)$.
	
	\item[(b)] The rate of convergence in (ii) depends on structural properties of the process (see, e.g., \cite{Douc-Fort-Guilin-2009} and \cite{Down-Meyn-Tweedie-1995}). 
	
	\item[(c)] The  rate of convergence  in (iii)  is necessarily exponential. Namely,  from (iii) we have that for any $t_0>0$,
	\begin{equation*}
	\lim_{n\to\infty}\sup_{x\in \R^d}\lVert p(nt_0,x,\D {y})
	-\uppi(\D {y})\rVert_{{\rm TV}} =0,
	\end{equation*} i.e., the Markov chain $\{X(nt_0)\}_{n\in\ZZ_+}$ is uniformly ergodic. Consequently, from \cite[Theorem 16.0.2]{Meyn-Tweedie-Book-2009} it follows that there exist $B,\beta>0$ such that 
	\begin{equation*}
	\lVert p(nt_0,x,\D {y})
	-\uppi(\D {y})\rVert_{{\rm TV}} \le B\E^{-\beta  n t_0},\qquad  x \in \R^d, \quad n\in\ZZ_+.\end{equation*} Since every $t\ge0$ can be clearly represented as $t=nt_0+s$ for some $n\in\ZZ_+$ and $s\in[0,t_0)$, we then conclude \begin{align*}\lVert p(t,x,\D {y})
	-\uppi(\D {y})\rVert_{{\rm TV}}&=\lVert p(nt_0+s,x,\D {y})
	-\uppi(\D {y})\rVert_{{\rm TV}}\\&\le\lVert p(nt_0,x,\D {y})
	-\uppi(\D {y})\rVert_{{\rm TV}}  \\&\le B\E^{-\beta  n t_0}
	\\&=B\E^{\beta s}\E^{-\beta  t}\\&\le B\E^{\beta t_0}\E^{-\beta  t},\qquad x \in \R^d, \quad t>0.\end{align*}
\end{enumerate}

\medskip

\noindent We are now in position to state the main result of this article.
Let $x_0\in\R^d$ and $r_0>0$ be as in (A4) and (A5). Set
\begin{align*}
A(x)&\df\frac{1}{2}{\rm Tr}\,\sigma(x)\sigma(x)^{T},\qquad x\in\R^{d},\\
B_{x_0}(x)&\df\langle x-x_0,b(x)\rangle,\qquad x\in\R^{d},\\
C_{x_0}(x)&\df\frac{\langle \sigma(x)^T(x-x_0),\sigma(x)^{T}(x-x_0)\rangle}{|x-x_0|^{2}},\qquad x\in\R^{d}\setminus\{x_0\},\\
\gamma_{x_0}(r)&\df\inf_{|x-x_0|=r}C_{x_0}(x),\qquad r\geq r_0,\\
\iota_{x_0}(r)&\df\sup_{|x-x_0|=r}\frac{2A(x)-C_{x_0}(x)+2B_{x_0}(x)}{C_{x_0}(x)},\qquad r\geq r_0,\\
I_{x_0}(r)&\df\int_{r_0}^{r}\frac{\iota_{x_0}(s)}{s}ds,\qquad r\geq r_0.
\end{align*}

\medskip

\begin{theorem}\label{tm:TV} Assume $\textbf{(A1)-(A5)}$, and  
	\begin{equation}\label{eq2}\Lambda\df\int_{r_0}^{\infty}\E^{-I_{x_0}(u)}\int_u^\infty  \frac{\E^{I_{x_0}(v)}}{\gamma_{x_0}(v)}\D v \,\D u<\infty\end{equation} (observe that (A5) ensures that $\Lambda$ is well defined).
	Then, $\{X(t)\}_{t\ge0}$ is uniformly ergodic.
\end{theorem}

\medskip

\noindent The   proof of Theorem \ref{tm:TV} is based on the Foster-Lyapunov method for uniform ergodicity of Markov processes, developed in \cite{Down-Meyn-Tweedie-1995}.
The method consists of identifying an appropriate recurrent (petite) set $C\in\mathcal{B}(\R^d)$ and constructing an appropriate  bounded function $L:\R^d\to[1,\infty)$ (the so-called Lyapunov  function) contained in the domain of  $\mathcal{G}$  (see \cite[Section 4]{Down-Meyn-Tweedie-1995} for details), such that the Lyapunov equation \begin{equation}
\label{eq:lyap}
\mathcal{G}L(x)\le-c_1L(x)+c_2\mathbb{1}_C(x),\qquad x\in\R^d,\end{equation} holds for some $c_1>0$ and $c_2\in\R$ (see \cite[Theorem 5.2]{Down-Meyn-Tweedie-1995}). 
Recall that    a set $C\in\mathcal{B}(\R^d)$ is said to be   petite if it satisfies the Harris-type minorization condition: there exist a probability   measure $\upeta_\mathcal{C}(\D t)$ on $\mathcal{B}((0,\infty))$ (the standard Borel $\sigma$-algebra on $(0,\infty)$) and a non-trivial measure $\upnu_C(\D x)$ on $\mathcal{B}(\R^d)$, such that $\int_0^\infty p(t,x,B)\upeta_C(\D t)\ge\upnu_C(B)$ for all $x\in C$ and $B\in\mathcal{B}(\R^d)$. Intuitively,  petite sets serve the role of singletons for Markov processes on non-discrete state spaces (see \cite[Chapter 5]{Meyn-Tweedie-Book-2009} and \cite[Section 4]{Meyn-Tweedie-AdvAP-II-1993}  for details).
However, the Lyapunov equation does not immediately imply ergodicity of $\{X(t)\}_{t\ge0}$. As in the discrete setting (see, e.g., \cite[Chapter 13]{Meyn-Tweedie-Book-2009}),  the process  has to satisfy additional structural properties (irreducibility and aperiodicity). Note that so far, we have not  used assumptions (A4) and (A5). In Proposition \ref{P1} we show that (A4) and (A5), together with (A1)-(A3), imply that  $\{X(t)\}_{t\ge0}$ is 

\medskip

\begin{enumerate}
	\item [(i)]	open-set irreducible, i.e., there exists a $\sigma$-finite measure $\uppsi(\D x)$ on
	$\mathcal{B}(\R^d)$ such that $${\rm supp}\,\uppsi\df\{x\in\R^d: \uppsi(O)>0\ \text{for every open neighborhood}\ O\ \text{of}\ x\}$$ has a non-empty interior, and whenever $\uppsi(B)>0$ we have
	$\int_0^{\infty}p(t,x,B)\D t>0$ for all $x\in\R^d$.
	
		\item [(ii)] aperiodic, i.e.,  
	there exist $t_0>0$ and a $\sigma$-finite measure $\upphi(\D x)$ on
	$\mathcal{B}(\R^d)$, such that $\upphi(B)>0$ implies
	$\sum_{n=0}^{\infty} p(nt_0,x,B) >0$ for all $x\in\R^d$.
\end{enumerate}

\medskip

\noindent Let us also remark that,  according to \cite[Propositio 4.1]{Meyn-Tweedie-AdvAP-II-1993} and  \cite[Theorems 5.1 and 7.1]{Tweedie-1994}, the $\mathcal{C}_b$-Feller property and open-set irreducibility of $\{X(t)\}_{t\ge0}$ ensure that the state space (in this case, $(\R^d,\mathcal{B}(\R^d)$) can be covered by a countable union of petite sets, and  that  every compact set is  petite.

To the best of our knowledge, Theorem \ref{tm:TV}  presents the first explicit conditions for the uniform ergodicity of diffusion processes based on the Lyapunov method. As previously noted, general conditions using the Lyapunov method for the uniform ergodicity of general Markov processes are derived in \cite{Down-Meyn-Tweedie-1995}. The only related results on this topic available in the literature (under the terminology of strong ergodicity) can be found in \cite[Section ]{Mao-2002}, \cite[Section 3]{Mao-2006} and \cite[Section 3]{Shao-Xi-2013}. The main difference is that the results in these works are based on the coupling method.   Therefore, the presented conditions are of the incremental type, which  in certain cases they can be  less practical, i.e. more difficult to verify (compare Examples \ref{ex1}, \ref{ex2} and \ref{ex3} with the examples in \cite{Mao-2002}, \cite{Mao-2006} and \cite{Shao-Xi-2013}).

\subsection{Literature review} \label{ss4}
Our work relates to active research on the ergodicity properties of Markov processes and the extensive literature on SDEs. 
In \cite{Arapostathis-Borkar-Ghosh_Book-2012}, \cite{Bhattacharya-1978}, \cite{Kulik-Book-2015}, \cite{Kulik-Book-2018}, \cite{Lazic-Sandric-2021}, \cite{Lazic-Sandric-2022}, \cite{Stramer-Tweedie-1997}, and \cite{Veretennikov-1997},   ergodicity properties of diffusion processes with respect to the total variation distance  are established using the Lyapunov(-type) method.
 In this article, building  on the ideas from \cite{Bhattacharya-1978} (see also \cite[Chapter 9]{Friedman-Book-1975} and \cite[Supplement]{Khasminskii-1960}), we complement these results by deriving  conditions 
for  the uniform ergodicity  of  these processes. Additionally, we apply these results to discuss the uniform ergodicity of a class of Markov jump processes obtained through the Bochner's subordination of diffusion processes. These results are related to 
\cite{Albeverio-Brzezniak-Wu-2010}, \cite{Arapostathis-Pang-Sandric-2019}, \cite{Deng-Schilling-Song-2017, Deng-Schilling-Song-Erratum-2018} \cite{Douc-Fort-Guilin-2009}, \cite{Down-Meyn-Tweedie-1995},
\cite{Fort-Roberts-2005}, \cite{Kevei-2018},  \cite{Kulik-2009},   \cite{Masuda-2007, Masuda-Erratum-2009}, \cite{Meyn-Tweedie-AdvAP-II-1993}, \cite{Meyn-Tweedie-AdvAP-III-1993}, \cite{Sandric-ESAIM-2016},  \cite{FYWang-2011}, \cite{Wang-2008}, \cite{Wang-2011}   and \cite{Wee-1999},  
where the ergodicity properties  of  general 
Markov processes are again established using the  Foster-Lyapunov method.  Analogous results have also been obtained in the discrete-time setting, see   \cite{Durmus-Fort-Moulines-Soulier-2004}, \cite{Douc-Moulines-Priouret-Soulier_Book-2018},   \cite{Fort-Moulines-2003}, \cite{Kulik-Book-2015}, \cite{Kulik-Book-2018}, \cite{Meyn-Tweedie-Book-2009}, \cite{Veretennikov-1997}, \cite{Veretennikov-2000}, \cite{Tuominen-Tweedie-1994} and the references therein.

The studies on ergodicity properties with respect to the total variation distance assume that 
the Markov processes are irreducible and aperiodic. This is satisfied if the process does not exhibit singular behavior in its motion, i.e., its diffusion part is non-singular and/or its jump part shows sufficient jump activity. Together with the Foster-Lyapunov condition, which ensures controllability of modulated moment of return times
to a petite set,  
these conditions lead to the ergodic properties stated.
In this article, we also discuss and provide conditions for the irreducibility and aperiodicity of diffusion processes. There is a vast literature on these, and related questions such as the strong
Feller property and heat kernel estimates of Markov processes. 
In particular, we refer the readers to 
\cite{Arisawa-2009}, \cite{Chen-Chen-Wang-2020}, \cite{Chen-Hu-Xie-Zhang-2017}, \cite{Chen-Zhang-2016},  \cite{Chen-Zhang-2018}, \cite{Grzywny-Szczypkowski-2019}, \cite{Kim-Lee-2019}, \cite{Kim-Song-Vondracek-2018}, \cite{Knopova-Schilling-2012}, \cite{Knopova-Schilling-2013},
	\cite{Kolokoltsov-2000}, \cite{Kolokoltsov-Book-2011}, \cite{Kwon-Lee-1999}, \cite{Pang-Sandric-2016},
	\cite{Sandric-TAMS-2016}, and \cite{Stroock-1975}
for the case of a class of Markov L\'evy-type processes with bounded coefficients,
and to
\cite{Arapostathis-Pang-Sandric-2019}, \cite{Bass-Cranston-1986}, \cite{Ishikawa-2001},
	\cite{Knopova-Kulik-2014}, \cite{Lazic-Sandric-2021}, \cite{Masuda-2007}, \cite{Masuda-Erratum-2009}, 
	\cite{Picard-1996, Picard-Erratum-2010}, \cite{Sandric-Valentic-Wang-2021}, \cite{Stramer-Tweedie-1997}, and \cite{Xi-Zhu-2019}
for the case of a class of It\^{o} processes.

At the end, we  remark that for Markov processes that do not converge in total variation (e.g., due to a lack of irreducibility and/or aperiodicity), 
ergodic properties under  Wasserstein distances are studied, as these processes may converge weakly under certain conditions. For instance, see \cite{Bolley-Gentil-Guillin-2012}, 
\cite{Butkovsky-2014}, \cite{Durmus-Fort-Moulines-2016}, \cite{Eberle-2011}, \cite{Eberle-2015}, \cite{Hairer-Mattingly-Scheutzow-2011}, \cite{Luo-Wang-2016}, \cite{Majka-2017}, \cite{Renesse-Sturm-2005} and \cite{Wang-2016}. 

\section{Proof of Theorem  \ref{tm:TV}}\label{s2}
In this section, we prove Theorem  \ref{tm:TV}. First, we prove that assumptions (A4) and (A5), together with (A1)-(A3), imply the open-set irreducibility and aperiodicity of  $\{X(t)\}_{t\ge0}$.

\begin{proposition}\label{P1} Under assumptions (A1)-(A5), the process 
	$\{X(t)\}_{t\ge0}$ is open-set irreducible and aperiodic.
	\end{proposition}
\begin{proof}
	We first show that (A5), together with (A1)-(A3), implies that $$\mathbb{P}^x(\uptau_{B_{r_0}(x_0)}<\infty)>0, \qquad  x\in\R^d,$$ where $x_0\in\R^d$ and $r_0>0$ are given in (A5), and $ \uptau_{B_{r_0}(x_0)}$ is defied as $\uptau_{B_{r_0}(x_0)}\df\inf\{t\ge0:X(t)\in B_{r_0}(x_0)\}$.
		Let $0<\varepsilon<r_0$, and define  $$\bar{L}(r)\df\int_{r_0-\varepsilon}^{r}\E^{-I_{x_0}(u)}\D u,\qquad r\geq r_0-\varepsilon.$$ 
	For $r> r_0-\varepsilon$ we then have
	$$
	\bar{L}'(r)=\E^{-I_{x_0}(r)}>0\qquad \text{and}\qquad 
	\bar{L}''(r)=-\frac{\bar{L}'(r)}{r}\iota_{x_0}(r).
	$$		
	Further, let $L:\R^{d}\to[0,\infty)$, $L\in \mathcal{C}^2(\R^d)$, be such that $L(x)=\bar{L}(|x-x_0|)$ for $x\in B^c_{r_0}(x_0)$. For $x\in B^c_{r_0}(x_0)$, we then have
	\begin{align*}
	\mathcal{G}L(x)&=\frac{1}{2}C_{x_0}(x)\bar{L}''(|x-x_0|)+\frac{\bar{L}'(|x-x_0|)}{2|x-x_0|}(2A(x)-C_{x_0}(x)+2B_{x_0}(x))\\
	&= \frac{\bar{L}'(|x-x_0|)}{2|x-x_0|}\left(2A(x)-C_{x_0}(x)+2B_{x_0}(x)-C_{x_0}(x)\iota_{x_0}(|x-x_0|)\right),
	\end{align*}
	and thus $$	\mathcal{G}L(x)\le0.$$
	Further, for $n\in\N$,  define 
	$\tau_n\df\uptau_{B^c_n(x_0)}.$ Clearly, $\uptau_n$, $n\in\N$, are stopping times such that (due to non-explosivity of $\{X(t)\}_{t\ge0}$) we have $\uptau_n\to\infty$, $\Prob^x$-a.s.,  as $n\to\infty$ for all $x\in\R^d$.
	Hence, the processes (recall that $\mathcal{G}$ is the extended generator of $\{X(t)\}_{t\ge0}$) $$L(X(t\wedge\tau_n))-L(X(0))-\int_0^{t\wedge\uptau_n}\mathcal{G}L(X(s))\D s,\qquad t\geq0\quad n\in\N,$$ are $\Prob^x$-martingales. For $x\in B^c_{r_0}(x_0)$, it then holds that 
	\begin{align*}
	\mathbb{E}^x[L(X(t\wedge\tau_n\wedge\tau_{B_{r_0}(x_0)}))]-\mathbb{E}^x[L(X(0))]&=\mathbb{E}^x\int_0^{t\wedge\tau_n\wedge\tau_{B_{r_0}(x_0)}}\mathcal{G}L(X(s))\D s\leq 0,
	\end{align*}
	i.e.,
	$$\mathbb{E}^x[\bar{L}(|X(t\wedge\tau_n\wedge\tau_{B_{r_0}(x_0)})-x_0|)]\le \bar{L}(|x-x_0|).$$
	In particular,
	$$\mathbb{E}^x[\bar{L}(|X(t\wedge\tau_n)-x_0|)\mathbb{1}_{\{\tau_{B_{r_0}(x_0)}>\tau_n\}}]\le \bar{L}(|x-x_0|),\qquad x\in B^c_{r_0}(x_0).$$ By letting $t\to\infty$, Fatou's lemma implies 
	$$\bar{L}(n)\mathbb{P}^x(\tau_{B_{r_0}(x_0)}>\tau_n)\le \bar{L}(|x-x_0|),\qquad x\in B^c_{r_0}(x_0).$$ Hence, by letting $n\to\infty$, we conclude that
	$$ \mathbb{P}^x(\tau_{B_{r_0}(x_0)}=\infty)\le \frac{\bar{L}(|x-x_0|)}{\bar L(\infty)}<1,\qquad x\in B^c_{r_0}(x_0),$$ i.e., $\mathbb{P}^x(\tau_{B_{r_0}(x_0)}<\infty)>0$ for all $x\in\R^d.$

	We next show that $\{X(t)\}_{t\ge0}$ is open-set irreducible. According  to \cite[Theorems 7.3.6 and 7.3.7]{Durrett-Book-1996}, there exists a strictly positive function  $q(t,x,y)$ on $(0,\infty)\times \bar{B}_{r_0}(x_0)\times\bar{B}_{r_0}(x_0)$, jointly continuous in $t$, $x$ and $y$, and twice continuously differentiable in $x\in B_{r_0}(x_0)$, satisfying  \begin{equation*} \mathbb{E}^x[f(X(t)) \mathbb{1}_{\{\uptau_{\bar{B}^c_{r_0}(x_0)}>t\}}]=\int_{B_{r_0}(x_0)}q(t,x,y)f(y)\D y,\end{equation*}   for any $t>0$, $x\in B_{r_0}(x_0)$ and any continuous, bounded  function $f:\R^d\to\R.$ By employing the dominated convergence theorem, the above relation holds also for $\mathbb{1}_O$, where $O\subseteq B_{r_0}(x_0)$ is an open set. Denote by $\mathcal{D}$ the class of all  $B\in\mathcal{B}(B_{r_0}(x_0))$ (the Borel $\sigma$-algebra on $B_{r_0}(x_0)$) such that \begin{equation*} \mathbb{P}^x(X_t\in B,\ \uptau_{\bar{B}^c_{r_0}(x_0)}>t)=\int_{B}q(t,x,y) \D y,\qquad t>0,\quad x\in B_{r_0}(x_0).\end{equation*} Clearly, $\mathcal{D}$ contains the $\pi$-system of open rectangles in $B_{r_0}(x_0)$, and forms a $\lambda$-system. By  Dynkin's $\pi$-$\lambda$ theorem, we conclude that $\mathcal{D}=\mathcal{B}(B_{r_0}(x_0)).$ Consequently, for any $t>0$, $x\in B_{r_0}(x_0)$ and $B\in\mathcal{B}(\R^d)$, we have  \begin{equation*} p(t,x,B)\ge \int_{B\cap B_{r_0}(x_0)}q(t,x,y)\D y.\end{equation*} Define  $\uppsi(\cdot)\df\uplambda(\cdot\cap B_{r_0}(x_0))$, where $\uplambda(\D x)$ denotes the Lebesgue measure on $\R^d$. Clearly, $\uppsi(\D x)$ is a $\sigma$-finite measure whose support has a non-empty interior. It remains to prove that  $$\int_0^{\infty}p(t,x,B)\D t>0,\qquad x\in\R^d,$$ whenever $\uppsi(B)>0$, $B\in \mathcal{B}(\R^d)$.
	For $x\in B_{r_0}(x_0)$ the implication holds trivially.
	Let  $x\in B^c_{r_0}(x_0)$  and $B\in\mathcal{B}(\R^d)$, $\uppsi(B)>0$, be arbitrary. For all $s>0$ we have
	\begin{align*}
	\int_0^\infty p(t,x,B) \D t & \ge \int_s^\infty  p(t,x,B)\D t \\&= \int_s^\infty \int_{\R^d} p(t-s,x,\D y)p(s,y,B) \D t \\
	&\geq \int_s^\infty \int_{B_{r_0}(x_0)} p(t-s,x,\D y)p(s,y,B) \D t \\
	&= \int_{B_{r_0}(x_0)}p(s,y,B) \int_s^\infty p(t-s,x,\D y) \D t .
	\end{align*}
	The assertion follows from the fact that $p(s,y,B)>0$ for $y\in B_{r_0}(x_0)$, and 
	\begin{equation*} \int_s^\infty p(t-s,x,B_{r_0}(x_0))\D t
	=\int_0^\infty p(t,x,B_{r_0}(x_0)) \D t
	=\mathbb{E}^x\left[\int_0^\infty \mathbb{1}_{\{X(t)\in B_{r_0}(x_0)\}}\D t\right]>0,
	\end{equation*}
	since $\{X(t)\}_{t\ge0}$ has continuous sample paths, $B_{r_0}(x_0)$ is an open set, and $\mathbb{P}^x(\uptau_{B_{r_0}(x_0)}<\infty)>0$ for every $x\in\R^d$.
	
	Finally, we show that $\{X(t)\}_{t\ge0}$ is aperiodic.  Specifically, we show that $$ \sum_{n=1}^\infty p(n,x,B)>0,\qquad x\in\R^d,$$ whenever $\uppsi(B)>0,$  where $B\in \mathcal{B}(\R^d)$.	For $x\in B_{r_0}(x_0)$, this relation clearly holds. For $x\in B^c_{r_0}(x_0)$ and $B\in\mathcal{B}(\R^d)$, with $\uppsi(B)>0$, we have \begin{equation*}
	\sum_{n=1}^\infty p(n,x,B)\ge\int_{B_{r_0}(x_0)}\sum_{n=1}^\infty p(n-t,x,\D y)p(t,y,B),\qquad t\in(0,1).
	\end{equation*} Since $p(t,y,B)>0$ for $y\in B_{r_0}(x_0)$, it suffices to show that \begin{equation*}\sum_{n=1}^\infty p(n-t,x,B_{r_0}(x_0))\ge\mathbb{P}^x\left(\bigcup_{n=1}^\infty\{X(n-t)\in B_{r_0}(x_0)\}\right)>0\end{equation*} for some $t\in(0,1)$. Now, assume by contradiction that this is not the case, i.e., \begin{equation*} \mathbb{P}^x\left(\bigcup_{n=1}^\infty\{X(n-t)\in B_{r_0}(x_0)\}\right)=0,\qquad t\in(0,1)\,.\end{equation*} This implies that \begin{equation*}\mathbb{P}^x\left(\bigcup_{q\in\mathbb{Q}_+\setminus\ZZ_+}\{X(q)\in B_{r_0}(x_0)\}\right)=0,\end{equation*} which is impossible, since $\{X(t)\}_{t\ge0}$ has continuous sample paths, $B_{r_0}(x_0)$ is an open set and $\mathbb{P}^x(\uptau_{B_{r_0}(x_0)}<\infty)>0$ for every $x\in\R^d$. Therefore, the original assumption must be false, and we conclude that $$ \sum_{n=1}^\infty p(n,x,B)>0$$ for all $x\in\R^d$ and $B\in \mathcal{B}(\R^d)$ with $\uppsi(B)>0,$  proving aperiodicity.
	\end{proof}

We are now in position to prove Theorem \ref{tm:TV}. To do so, we  apply \cite[Theorem 5.2]{Down-Meyn-Tweedie-1995}. 
However, note that the notion of aperiodicity used in \cite[Theorem 5.2]{Down-Meyn-Tweedie-1995} differs slightly from the one used in this article. Therefore, we need to show that the  aperiodicity defined in this article implies the notion used in \cite{Down-Meyn-Tweedie-1995}. 
Indeed, by
\cite[Proposition 6.1]{Meyn-Tweedie-AdvAP-II-1993}, \cite[Theorem 4.2]{Meyn-Tweedie-AdvAP-III-1993}, along with the 
aperiodicity of $\{X(t)\}_{t\ge0}$, there exists  a petite set $C\in\mathcal{B}(\R^d)$, a time 	 $T>0$, and a non-trivial measure $\upnu_C(\D x)$ on $\mathcal{B}(\R^d)$, such that $\upnu_C(C)>0$ and $$p(t,x,B)\ge\upnu_C(B),\qquad x\in C,\quad t\ge T,\quad B\in\mathcal{B}(\R^d).$$ In particular, we have
$$p(t,x,C)>0,\qquad x\in C,\quad t\ge T,$$	
which aligns exactly with the definition of aperiodicity used in \cite{Down-Meyn-Tweedie-1995}.

\begin{proof}[Proof of Theorem \ref{tm:TV}]
Define
$$\bar{L}(r)\df\int_{r_0}^{r}\E^{-I_{x_0}(u)}\int_u^{\infty}\frac{\E^{I_{x_0}(v)}}{\gamma_{x_0}(v)}\D v\,\D u,\qquad r\geq r_0.$$  Clearly, for  $r\ge r_0$ it holds that $\bar{L}(r)\le\Lambda$,
 \begin{align*}\bar{L}'(r)=\E^{-I_{x_0}(r)}\int_r^{\infty}\frac{\E^{I_{x_0}(u)}}{\gamma_{x_0}(u)}\D u\qquad\text{and}\qquad
\bar{L}''(r)=-\iota_{x_0}(r)\frac{\bar{L}'(r)}{r}-\frac{1}{\gamma_{x_0}(r)}.
\end{align*}
Further, fix $r_1>r_0$  and let $L:\R^{d}\to[0,\infty)$, $L\in \mathcal{C}^2(\R^d)$, be such that $L(x)=\bar{L}(|x-x_0|)+1$ for $x\in B^c_{r_1}(x_0).$ Now, for $x\in B^c_{r_1}(x_0)$, we have (recall the representation of $\mathcal{G}$ given in \eqref{gen})
\begin{align*}
\mathcal{G}L(x)&=\frac{1}{2}C_{x_0}(x)\bar{L}''(|x-x_0|)+\frac{\bar{L}'(|x-x_0|)}{2|x-x_0|}(2A(x)-C_{x_0}(x)+2B_{x_0}(x))\\
&\le \frac{\bar{L}'(|x-x_0|)}{2|x-x_0|}(2A(x)-C_{x_0}(x)+2B_{x_0}(x)-C_{x_0}(x)\iota_{x_0}(x))-\frac{C_{x_0}(x)}{2\gamma_{x_0}(|x-x_0|)}
\\
&\le-\frac{C_{x_0}(x)}{2\gamma_{x_0}(|x-x_0|)}\\
&\le -\frac{1}{2}\\
&\le-\frac{1}{2(\Lambda +1)}L(x).
\end{align*}
 Thus, we have obtained the relation in \eqref{eq:lyap} with $c_1=1/(2(\Lambda +1))$, $C=\bar B_{r_1}(x_0)$ and $c_2=\sup_{x\in C}|\mathcal{G}L(x)|$, which proves the assertion of the theorem.
 \end{proof}

 \section{Examples and Applications of  Theorem  \ref{tm:TV}}\label{s3}
 In this section, we provide several examples of diffusion processes that satisfy the assumptions of  Theorem  \ref{tm:TV}, along with an application of Theorem  \ref{tm:TV} to a class of Markov processes with jumps.
We begin with a simple example of a diffusion process with polynomial drift.
 
 \begin{example}\label{ex1} Let $\sigma(x)=I_d$ (the $d\times d$ identity matrix) and let $b(x)=-Kx|x|^{\kappa-1}$ with $K>0$ and $\kappa>1$. It is straightforward to verify that $b(x)$ and $\sigma(x)$ satisfy assumptions (A1)-(A5) with $x_0=0$ and arbitrary $r_0>0$, and  that $\gamma_{x_0}(r)=1$ and $$I_{x_0}(r)=\frac{2K}{\kappa+1}r_0^{\kappa+1}-\frac{2K}{\kappa+1}r^{\kappa+1}.$$
 	We then have $$\Lambda=\int_{r_0}^\infty\E^{\frac{2K}{\kappa+1}u^{\kappa+1}}\int_u^\infty\E^{-\frac{2K}{\kappa+1}v^{\kappa+1}}\D v\,\D u.$$
 	Observe that \begin{align*}\int_u^\infty\E^{-\frac{2K}{\kappa+1}v^{\kappa+1}}\D v&=(2K)^{\frac{-2\kappa-1}{\kappa+1}}(\kappa+1)^{\frac{\kappa}{\kappa+1}}\int_{\frac{2K}{\kappa+1}u^{\kappa+1}}v^{-\frac{\kappa}{\kappa+1}}\E^{-v}\D v\\
 	&=(2K)^{\frac{1}{\kappa}}(\kappa+1)^{-\frac{\kappa+1}{\kappa}}\Gamma\left(\frac{1}{\kappa+1},\frac{2K}{\kappa+1}u^{\kappa+1}\right),\end{align*}
where $\Gamma(p,q)$, $p,q\in\R$, is the incomplete gamma function. According to \cite[display 6.5.32]{Abramowitz-Stegun-1964}, $$\lim_{u\to\infty}\frac{\Gamma\left(\frac{1}{\kappa+1},\frac{2K}{\kappa+1}u^{\kappa+1}\right)}{u^{-\kappa}\E^{-\frac{2K}{\kappa+1}u^{\kappa+1}}}=2^{-\frac{\kappa}{\kappa+1}}(\kappa+1)^{\frac{\kappa}{\kappa+1}},$$
 which proves that $\Lambda<\infty.$ \qed
 	\end{example}
 
In the next example, we consider a one-dimensional diffusion process with a non-incremental drift coefficient.
 
 \begin{example}\label{ex2}
	Let $\sigma(x)=1$,  and let $b(x)$ be locally Lipschitz continuous, such that $b(x)=-Kx|x|^{\kappa-1}(\cos x+\rho)$ for $x\in B^c_1(0)$, where $K>0$, $\kappa>1$, and $\rho>0$. It is straightforward to verify that $b(x)$ and $\sigma(x)$ satisfy assumptions (A1)-(A5) with $x_0=0$ and arbitrary $r_0>0$. Furthermore,  we have   $\gamma_{x_0}(r)=1$ and $$I_{x_0}(r)=\frac{2K\rho}{\kappa+1}r_0^{\kappa+1}-\frac{2K\rho}{\kappa+1}r^{\kappa+1}-2K\int_{r_0}^ru^{\kappa}\cos u\,\D u.$$
	Note that $$\lim_{r\to\infty}\frac{\int_{r_0}^ru^{\kappa}\cos u\,\D u}{r^{k+\varepsilon}}=\lim_{r\to\infty}\frac{r^k\sin r-r_0^k\sin r_0-\kappa\int_{r_0}^ru^{\kappa-1}\sin u\,\D u}{r^{k+\varepsilon}}=0$$ for any $\varepsilon>0$. Thus, the finiteness of $\Lambda$ reduces on showing the finiteness of $$\int_{r_0}^\infty\E^{\frac{2K\rho}{\kappa+1}u^{\kappa+1}}\int_u^\infty\E^{-\frac{2K\rho}{\kappa+1}v^{\kappa+1}}\D v\,\D u,$$ which, as  shown in Example \ref{eq1}, is finite if, and only if, $\kappa>1$.
	\qed
\end{example}

 In the following example, we discuss uniform ergodicity of a class of Langevin tempered diffusion processes.

  \begin{example}\label{ex3}
 	Let $\alpha\in(0,1/d)$, and let $\pi\in \mathcal{C}^2(\R^d)$ be strictly positive, defined as 
 	$\pi(x)\df c|x|^{-\nicefrac{1}{\alpha}}$ for some $c>0$ and for all $x\in B_1^c(0)$. Additionally, assume that
 	 $\int_{\R^d}\pi(x)\D x=1$.
 	For $\beta\in(0,(1+\alpha(2-d))/2)$ and $x\in\R^d$,  define
 	\begin{equation*}
 	\sigma(x)\df(\pi(x))^{-\beta}I_d,
 	\end{equation*}
 	and
 	\begin{equation*}
 	b(x)\df\frac{1}{2}(\sigma(x)\sigma(x)^T\nabla\log(\pi(x))+\nabla\cdot \sigma(x)^T\sigma(x))
 	=\frac{1-2\beta}{2}\pi(x)^{-2\beta}\nabla\log(\pi(x)).
 	\end{equation*}
 	Observe that $$\sigma(x)=c^{-\beta}|x|^{\frac{\beta}{\alpha}}\quad\text{and}\quad b(x)= -\frac{1-2\beta}{2\alpha}c^{-2\beta}x|x|^{\frac{2\beta}{\alpha}-2},\qquad x\in B_1^c(0).$$
 Note  that $b(x)$ and $\sigma(x)$ do not in general satisfy assumptions (A2) and A(3). However, observe that in the proof of Theorem \ref{tm:TV} we  did not actually use the fact that $\{X(t)\}_{t\ge0}$ is a unique strong solution to \eqref{eq1}. All that we required is that the martingale problem for $(b,\sigma\sigma^T)$ is well posed, which is equivalent to  \eqref{eq1} admitting a unique (in distribution) weak solution (see \cite[Theorem V.20.1]{Rogers-Williams-Book-II-2000}). Furthermore,  in \cite[Proposition 15]{Fort-Roberts-2005} it has been shown that the SDE in \eqref{eq1}
 admits a weak solution
 $(\Omega,\mathcal{F},\{\mathcal{F}(t)\}_{t\ge0},\{B(t)\}_{t\ge0},\{X(t)\}_{t\ge0},\Prob)$,
which is a conservative strong Markov process with continuous sample paths.
Moreover, it is open-set irreducible, aperiodic, every compact set is petite, and
$\uppi(\D x)\df\pi(x)\D x$ is its unique invariant probability measure.
Here, $(\Omega,\mathcal{F}, \{\mathcal{F}(t)\}_{t\ge0},\{B(t)\}_{t\ge0},\Prob)$ is a
standard $n$-dimensional Brownian motion.
 Now, take $x_0=0$ and $r_0>1$.	It is straightforward to verify   that $\gamma_{x_0}(r)=c^{-2\beta}r^{\frac{2\beta}{\alpha}}$ and $$I_{x_0}(r)=\frac{1-2\beta}{\alpha}\ln r_0-\frac{1-2\beta}{\alpha}\ln r.$$
 		We then have $$\Lambda=c^{2\beta}\int_{r_0}^\infty\E^{\frac{1-2\beta}{\alpha}\ln u}\int_u^\infty\frac{\E^{-\frac{1-2\beta}{\alpha}\ln v}}{v^{\frac{2\beta}{\alpha}}}\D v\,\D u=c^{2\beta}\int_{r_0}^\infty u^{\frac{1-2\beta}{\alpha}}\int_u^\infty v^{-\frac{1}{\alpha}}\D v\,\D u.$$ This simplifies to $$\Lambda=\frac{\alpha c^{2\beta}}{1-\alpha}\int_{r_0}^\infty u^{1-\frac{2\beta}{\alpha}}\D u,$$ which is finite if, and only if,  $\alpha<\beta.$\qed
 	 \end{example}

The results discussed in this article can be  used to establish uniform ergodicity of a class of Markov processes with jumps. Recall that a subordinator  $\{S(t)\}_{t\ge0}$ is a non-decreasing L\'{e}vy process on $\left[0,\infty\right)$. If $\{X(t)\}_{t\ge0}$ and $\{
 S(t)\}_{t\ge0}$ are independent, 
 then the process $X^{S}(t)\df X(S(t))$, $t\ge0$, obtained by applying random time change to $\{X(t)\}_{t\ge0}$   through $\{S(t)\}_{t\ge0}$, is referred to as the subordinate process $\{X(t)\}_{t\ge0}$ with subordinator $\{S(t)\}_{t\ge0}$ in the sense of Bochner. It is easy to see that $\{X^S(t)\}_{t\ge0}$ is  a Markov process with  the transition kernel
 $$p^S(t,x,\D y)=\int_{\left[0,\infty\right)} p(s,x,\D y)\mu_t(\D s),$$
 where $\upmu_t(\D s)=\mathbb{P}(S(t)\in\D s)$ is the transition probability of $S(t)$, $t\ge0$. Additionally, it is straightforward to check that if $\uppi(\D x)$ is an invariant probability measure for $\{X(t)\}_{t\ge0}$, then it is also invariant for the subordinate process $\{X^S(t)\}_{t\ge0}$.
 Moreover, if $$\lVert p(t,x,\D {y})
 -\uppi(\D {y})\rVert_{{\rm TV}}\le B(x)r(t),\qquad x\in\R^d,\quad t\ge0,$$ for some $B:\R^d\to[0,\infty)$ and measurable $r:[0,\infty)\to[0,\infty)$,
 then
 \begin{align*}
 \lVert p^S(t,x,\D {y})
 -\uppi(\D {y})\rVert_{{\rm TV}}&=\left\lVert \int_{\left[0,\infty\right)} (p(s,x,\D y)-\uppi(\D {y}))\upmu_t(\D s)
 \right\rVert_{{\rm TV}}\\&\le c(x)\int_{\left[0,\infty\right)}r(s)\upmu_t(\D s)\\
 &=B(x)\mathbb{E}[r(S(t))].
 \end{align*}
 Furthermore, since $\mathbb{E}[S(t)]>0$ for $t>0$, the strong law of large numbers (see, e.g., \cite[Theorems 36.5 and 36.6]{Sato-Book-1999}) implies that $\lim_{t\to\infty} S(t)=\infty$, almost surely. Hence, if $r(t)=\E^{-\beta t}$ for some $\beta>0$, the dominated convergence theorem gives that $$\lim_{t\to\infty}\mathbb{E}[r(S(t))]=0.$$ Therefore, if  $\{X(t)\}_{t\ge0}$ is uniformly ergodic, then $\{X^S(t)\}_{t\ge0}$ is also uniformly ergodic for any subordinator $\{S(t)\}_{t\ge0}$.

\section*{Acknowledgement} 
\noindent The author thanks the anonymous referee for the helpful comments that have led to significant
improvements of the results in the article.
This work has been supported by \textit{Croatian Science Foundation} under project 2277.

\section*{Conflict of interest}
 \noindent The author declares that he has no conflict of interest.

\bibliographystyle{abbrv}
\bibliography{References}

\end{document}